# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v6

@@ Abstract
Index Terms:
ADAS, intelligent vehicle, foggy lane detection dataset, foggy lane detection, global attention, edge enhancement
I Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[1].
@@ I Introduction
Figure 1: Lane detection in foggy scenarios for ADAS.[2]
@@
Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting. Son et al.[3] first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.[4] used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
@@
With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings[5, 6, 7, 8, 9, 10]. However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.[11] addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
@@
In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection[12, 13, 14, 15]. Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable. In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position[16, 17, 18, 19, 20, 21, 22]. These innovations improve accuracy and efficiency in complex or occluded scenarios.
@@
Research on downstream tasks in adverse weather conditions[23], such as fog, is limited compared to other detection and segmentation tasks. Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results. Some methods combine image enhancement with detection to reduce weather interference[24], while others simulate foggy images to augment datasets. For example, Nie et al.[25] generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
@@
Domain adaptation techniques also address the style gap between foggy and clear-weather images[26, 27]. Hnewa et al.[28] proposed a multi-scale domain-adaptive YOLO model, while Li et al.[29] introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
@@
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. In addition, we generate foggy versions of two widely-used lane detection datasets, CULane [8] and Tusimple [30], creating FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@ II Related Work
II-A Segmentation-based Lane Detection
@@
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet[5] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN[8] extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA[31] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer[32] presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
II-B Curve-based Lane Detection
@@
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet[33] introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR[34] can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet[35] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection. PGA-Net[36] introduces an end-to-end framework for lane detection, utilizing cubic polynomial functions and a transformer-based DETR model to simultaneously predict lane shape parameters and incorporate global road context, with the addition of Mean Curvature Loss to enhance the accuracy of curve predictions.
II-C Anchor-based Lane Detection
@@
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN[37] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT[17] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet[18] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet[19] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
II-D Row-wise-based Lane Detection
@@
Row-wise methods predict lane positions for each image row. E2E-LMD[38] first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions. UFLD[11] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet[39] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
III Data Constrution
@@
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [8] and Tusimple [30] datasets to test the generalizability of our methods.
@@
Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[40]
@@
Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[40]
III-A FoggyLane
@@
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane[8], Tusimple[30], and LLAMAS[41], lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
@@
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
@@
Comparing FoggyLane with existing datasets, such as Caltech-Lanes[43] and Tusimple[30], which focus on single traffic scenarios, and LLAMAS[41], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape [44] and BDD100k [45] include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane [8] and CurveLanes [46] cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple [30] and CULane [8] formats for broader accessibility.
@@
TABLE I: Comparison Between Our Dataset and Existing Lane Detection Datasets
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [43] 1.2k None 640 x 480 Tusimple [30] 6.4k None 1280 x 720 LLAMAS [41] 100k None 1276 x 717 Appolloscape [44] 110k None 3384 x 2710 BDD100k [45] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [46] 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
@@
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[47]. These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
@@ III-B FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[30]. The CULane dataset [8] is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing. In contrast, the TuSimple dataset [30] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model [25], and the specific generation method is provided in the Supplementary Material for Review. As shown in Figure 5, it demonstrates the generation case of FoggyCULane.
IV Methods
IV-A Overall structure

### v7

@@ Abstract
keywords:
[…]
1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Olovsson et al., 2022).
@@ 1 Introduction
Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2). This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Lim et al., 2021). In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments. Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions. Research related to vehicle trajectory analysis (Berghaus et al., 2024), congestion-aware traffic prediction (Zhang et al., 2025b), and 3D scene understanding (Qu et al., 2025) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels, 2025)
@@
Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting. Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
@@
With deep learning’s success in image tasks, it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018). However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
@@
In addition to segmentation methods, lane detection also includes detection-based approaches, many borrowing Anchor-based methods from object detection(Redmon, 2016; Farhadi and Redmon, 2018). Due to the elongated shape of lane lines, traditional bounding boxes are unsuitable. In lane detection, anchors are often preset starting points and lines with different slopes, with the network predicting the relative distance between the lane and the anchor to optimize shape and position(Chen et al., 2019; Zheng et al., 2022; Ran et al., 2023; Wang et al., 2025). These innovations improve accuracy and efficiency in complex or occluded scenarios.
@@
Research on downstream tasks in adverse weather conditions, such as fog, is limited compared to other detection and segmentation tasks. Traditional solutions often use image enhancement techniques, like dehazing, to improve visual clarity, but this doesn’t always lead to better detection results. Some methods combine image enhancement with detection to reduce weather interference(Huang et al., 2020), while others simulate foggy images to augment datasets. For example, Nie et al.(Nie et al., 2022) generated foggy images from clear lane images in the CULane dataset, achieving synthetic data augmentation that improved recognition rates from 76.65% to 86.65% under light fog.
@@
Domain adaptation techniques also address the style gap between foggy and clear-weather images(Chen et al., 2018; Zhang et al., 2021). Hnewa et al.(Hnewa and Radha, 2021) proposed a multi-scale domain-adaptive YOLO model, while Li et al.(Li et al., 2023) introduced a transfer learning framework for fog-adaptive object detection using labeled clear-weather and unlabeled foggy images, avoiding the need for annotated foggy datasets.
@@
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. In addition, we generate foggy versions of two widely-used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), creating FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@ 2 Related Work
2.1 Segmentation-based Lane Detection
@@
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN(Pan et al., 2018) extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
2.2 Curve-based Lane Detection
@@
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet(Tabelini et al., 2021b) introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR(Liu et al., 2021b) can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
2.3 Anchor-based Lane Detection
@@
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN(Li et al., 2019) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT(Tabelini et al., 2021a) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet(Zheng et al., 2022) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet(Ran et al., 2023) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
2.4 Row-wise-based Lane Detection
@@
Row-wise methods predict lane positions for each image row. E2E-LMD(Yoo et al., 2020) first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions. UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
3 Data Constrution
@@
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
@@
Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.(GoogleMaps, Accessed: 2024)
@@
Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.(GoogleMaps, Accessed: 2024)
3.1 FoggyLane
@@
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane(Pan et al., 2018), Tusimple(Tusimple, 2017), and LLAMAS(Behrendt and Soussan, 2019), lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
@@
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
@@
Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
@@
Table 1: Comparison Between Our Dataset and Existing Lane Detection Datasets
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
@@
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.(Zhang et al., 2024a). These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
Figure 5: Networks of foggy image generation. $\otimes$ represents element-wise multiplication.
@@ 3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017). The CULane dataset (Pan et al., 2018) is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing. In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model (Nie et al., 2022), and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder (Harald, 1924) is represented as:
$I(x)=J(x)t(x)+A(1-t(x))$ (1)
where $x$ indicates a certain pixel of the image, $J(x)$ refers to the original clear image of the object and $I(x)$ refers to the foggy image. $A$ is the atmospheric light value at infinity and $t(x)$ can be further defined as:
$t(x)=e^{-\beta d(x)}$ (2)
where $\beta$ is the extinction coefficient, which is closely related to the concentration, size, type and distribution of particulate matter. $d(x)$ represents the distance between the object and the viewer. As this distance increases, the transmittance decreases, resulting in thicker fog.
Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information. The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image. After passing through the DSA Module and STE Module, as shown in Fig. 5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map. The Depth Prediction (DP) is responsible for depth estimation.
According to the transmittance formula Eq. (2), we use the dark channel prior(He et al., 2011) to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
$I_{\text{dark }}(x)=\min_{c\in R,G,B}\left(\min_{y\in\Omega(x)}I_{c}(y)\right)$ (3)
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 6 and 7.
4 Methods
4.1 Overall structure